OpenStudy (anonymous):

Find the slope of a line parallel to 5x+2y=6 A- 2/5 B- -2/5 C- 5/2 D- -5/2

OpenStudy (anonymous):

either a) solve for \(y\) or b) memorize that the slope of \(ax+by=c\) is \(-\frac{a}{b}\)

OpenStudy (anonymous):

in your case \(a=5,b=2\)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got this?

OpenStudy (anonymous):

another way would be to rearrange it into point slope form so basically solve for y 5x+2y=6 -5x -5x 2y=6-5x divide both sides by two y=3-5/2x or y=-5/2x+3 and you have it in y=mx+b from and remember m=slope